OpenStudy (anonymous):

Identify the slope and y-intercept of the inequality 4x + 2y > 8.

OpenStudy (anonymous):

slope = -2, y-intercept = 4 slope = 2, y-intercept = 4 slope = -2, y-intercept = 8 slope = - 1/2 , y-intercept = 4

OpenStudy (anonymous):

ALright, first solve for y. 4x + 2y > 8 2y > -4x + 8 y > -2x + 4 This is the slope-intercept form. y = mx + b m = slope b = y-intercept

OpenStudy (anonymous):

So you understand now?

OpenStudy (anonymous):

Sorry I stepped away for a minute.

OpenStudy (anonymous):

-2 would be the slope, and 4 would be the intercept?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yay!, thanks (:

OpenStudy (anonymous):

You're welcome :)